OpenStudy (anonymous):

Simplify the expression 5x − (3x2 + 9x).

OpenStudy (callisto):

5x − (3x^2 + 9x) = 5x -3x^2 - 9x = -3x^2 + 5x - 9x = -3x^2 - 4x